# Giuseppe Scanniello+

### v1

@@ Abstract
The management of chronic Heart Failure (HF) presents significant challenges in modern healthcare, requiring continuous monitoring, early detection of exacerbations, and personalized treatment strategies. In this paper, we present a predictive model founded on Machine Learning (ML) techniques to identify patients at HF risk. This model is an ensemble learning approach, a modified stacking technique, that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models. We initially assess the model on a real dataset and the obtained results suggest that it performs well in the stratification of patients at HR risk. Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified. As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts. However, it is acceptable given our priority of identifying patients at risk of HF because they will be asked to participate in the telemonitoring program of the PrediHealth research project on which some of the authors of this paper are working. The initial findings also suggest that ML-based risk stratification models can serve as valuable decision-support tools not only in the PrediHealth project but also for healthcare professionals, aiding in early intervention and personalized patient management. To have a better understanding of the value and of potentiality of our predictive model, we also contrasted its results with those obtained by using three baseline models. The preliminary results indicate that our predictive model outperforms these baselines that flatly consider features, i.e., not grouping them in clinical and echocardiographic features.
@@ I Introduction
The human heart works like a pump, sending blood and oxygen around the body. A major cause of morbidity and mortality worldwide is chronic Heart Failure (HF), which is defined by the heart’s inability to pump blood efficiently. Lifestyle risk factors such as smoking, poor diet, and high blood pressure are closely associated with the increase in the incidence of heart disease. Current techniques often fail to provide prompt and precise identification of heart disease, even with advances in medical diagnostics [2, 3]. It is also worth mentioning that cardiovascular diseases are associated with the highest number of deaths among Non-Communicable Diseases, amounting to 17.9 million deaths annually, followed by cancers (9.3 million), chronic respiratory diseases (4.1 million), and diabetes (2 million, including kidney disease linked to diabetes) [4].
@@
The PrediHealth project aims to address the challenges above by integrating telemedicine, mobile health (m-health) solutions, and predictive analytics into an IoT healthcare platform. Specifically, it leverages an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and ML/AI-driven predictive models to support clinical decisions. As it is easy to follow, at the core of the PrediHealth project lies the predictive model–designed for the stratification of patients at risk of HF–that is presented in this paper for the first time. This model is based on an ensemble learning approach–a modified stacking technique–that uses two specialized models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these two models. In this paper, we detail the model’s methodological foundations and its initial empirical assessment, as well as the possible implications for our results outside the goal of the PrediHealth project. We assessed the validity of the model by conducting an initial empirical study. The obtained results suggest that the model performs pretty well. Specifically, we obtained high sensitivity (95%), ensuring that nearly all high-risk patients are identified. As for accuracy, we obtained 84%, which can be considered moderate in some ML contexts. However, it is acceptable in the PrediHealth project, given the priority of identifying patients at risk of HF to be then included in a telemonitoring program, to act promptly to let them receive medical attention, thus hopefully preventing them from dying. We also compared our model against three baseline models (Decision Tree, Random Forest, and Support Vector Regression). Our initial outcomes suggest that the use of our predictive model outperforms these baselines (e.g., in terms of the accuracy and the sensitivity of the patient stratification).
@@
Paper Structure. The following sections will explore the model’s design and evaluation strategies, providing a perspective on its role in the future of predictive healthcare and in PrediHealth, in particular. Precisely, the rest of the paper is organized as follows. In Section II, we introduce the PrediHealth project and present related research and background information, while the design of the predictive model is shown in Section III. The design of the study to empirically assess our predictive model and the results of this assessment are reported in Section IV and Section V, respectively. In this latter section, we also discuss the results and delineate possible implications for our research. We conclude the paper in Section VI with final remarks and future work.
@@ II Background and Related Work
In this section, we first provide information on the PrediHealth project. It is the project for which the predictive model presented in this paper has been developed. We conclude by motivating our work and discussing related research.
@@ II-A The PrediHealth Project
We restricted ourselves to a few aspects of the project for space reasons and scant relevance to the overarching goals of the research presented in the paper. Readers interested in further details can refer to .
@@ II-B Predictive Models in Digital Health
Other researchers have explored ML-driven models for HF prognosis, utilizing algorithms like logistic regression, decision trees, random forests, support vector machines (SVMs), and ensemble classifiers. For instance, Pinem et al. [6] demonstrated that a Voting Classifier combining Logistic Regression and SVM achieved the highest accuracy (88.04%) and ROC_AUC score (88%), with blood pressure and cholesterol emerging as key predictive factors. Similarly, Guo et al. [7] employed ML techniques to predict HF outcomes using EHR data, incorporating medical notes, laboratory results, and imaging data to achieve expert-level diagnostic accuracy. These studies underscore the role of ML in early HF detection, risk assessment, and personalized intervention strategies. Recently, Visco et al. [8] developed a predictive model for cardiovascular diagnostics. Their study identified correlations between worsening HF and three key parameters: Creatinine, sPAP, and CAD. To enhance interpretability, they introduced a GP-based classifier and extracted from it a formula that achieved almost 97% accuracy, outperforming traditional ML models. On the other hand, the authors themselves identified how the use of only four features, despite the high accuracy, could leave out from the doctor’s final evaluation other parameters that could be important in the final prediction.
@@
The research presented in this paper aims to bridge these gaps by developing a robust ML-based predictive model that integrates diverse patient datasets, including HFpEF cases, to enhance generalizability. The proposed model not only refines risk stratification but also enhances clinical applicability by ensuring interpretability and reliability across heterogeneous patient populations. In addition, our predictive model is integrated within advanced solutions in the context of CPS/IoT fields [^2] While CPS encompasses a broader range of networked systems, IoT specifically focuses on internet-enabled connectivity. by integrating state-of-the-art technologies such as telemonitoring and m-health to facilitate the integrated hospital-territory-home management of HF patients.
@@ III Design of the Predictive Model
In this section, we first introduce the dataset employed for training and testing the model. We then describe the architecture of the proposed model in detail.
@@ III-A Dataset
The used dataset contains real data. That is, it comprises clinical records of patients diagnosed with heart failure, collected as part of a structured medical registry. It encompasses a diverse set of 1040 patient records, each characterized by 66 features that provide a comprehensive view of patient demographics, clinical history, treatment details, and outcomes. The dataset includes key variables such as age, height, weight, and body mass index (BMI), along with clinical indicators like ejection fraction (EF), New York Heart Association (NYHA) classification, and specific diagnoses, including ischemic, hypertensive, and dilated cardiomyopathy. Additionally, dataset documents prescribed medications, including ACE inhibitors and aldosterone antagonists, offering insights into pharmacological management. Longitudinal follow-up data track patient progress, recording mortality outcomes such as cardiac death, heart failure exacerbation, sudden death, and interventions like heart transplants or ventricular assist device (VAD) implantation. The dataset also contains other bio-chemical pieces of information like Na+ (Sodium) or Uric Acid.
Before using the dataset, we preprocessed it to ensure data consistency and reliability. In other words, the dataset underwent pre-processing steps to address missing values, standardize feature formats, and encode categorical variables where necessary, ensuring consistency and reliability for subsequent modeling and evaluation. In the following sections, we detail the performed pre-processing steps and outline the criteria and methodology used for the selection of the features used in our predictive model, explaining the rationale behind the choice of these features and their relevance for the construction of such a model.
III-A1 Pre-processing
The dataset underwent the following pre-processing steps.
  • The gender encoding was standardized, as the dataset was originally formed by merging two datasets that used different encoding schemes. In one dataset, males were labeled as 1 and females as 0, whereas in the other, females were encoded as 1 and males as 2. To unify this, the gender values were mapped to a common scheme where male = 0 and female = 1.
  • All rows missing the target variable “death” were removed to maintain the integrity of the supervised learning process. Following this, rows containing more than three missing values were eliminated to reduce the impact of incomplete records on model performance.
  • Date values were converted into a standardized date-time format, and a new feature, “lifespan,” was introduced to represent the number of days between the initial characterization date and the recorded outcome date (“death”). Cases where the lifespan was negative were manually inspected and removed to ensure logical consistency in survival time calculations.
  • Outliers in the “lifespan” variable were identified and removed using the inter-quartile range (IQR) method. The lower and upper bounds were determined based on 1.5 times the IQR, and values outside this range were filtered out. The cleaned lifespan distribution was then analyzed to ensure a more balanced dataset.
  • As for the outcome variable (or also dependent variable), we used a binary classification for the values in the dataset. We employed a threshold-based labeling approach, using a 440-day threshold to classify patients into two groups: “at risk” and “not at risk.” Patients with a lifespan greater than or equal to 440 days were assigned label = 0, while those with a lifespan below this threshold were assigned label = 1. This transformation allowed for a more structured and interpretable analysis, enabling predictive modeling focused on early mortality risk assessment. The choice of this threshold is justified by technical reasons and clinical bases. Indeed, from a technical point of view, this threshold divides the dataset almost perfectly; in fact, we have 48% of samples with the label ”at risk” and 52% of samples with the label ”not at risk.” This also allowed the predictive model to exhibit good prediction performance of the outcome values. As far as the clinical base is concerned, the used threshold can be considered to be compliant with a period within which vigilant monitoring and proactive management are crucial to mitigate adverse outcomes in HF patients.
After completing all pre-processing steps shown just above, the final dataset consisted of 182 patient records to be used in our predictive model.
III-A2 Features
@@
In our predictive model, we selected 33 out of 66 available features, including clinical and echocardiographic parameters relevant to cardiac evaluation. The chosen characteristics encompass demographic attributes such as age, BMI and sex, along with essential clinical indicators such as ejection fraction (EF), NYHA classification, NT-proBNP, creatinine, glucose, and hemoglobin (Hb). Additionally, the dataset contains information on comorbidities such as hypertension, diabetes, dyslipidemia, and respiratory diseases. The selection of features was based on the experience some of the authors gained in the field of HF, taking into account the purposes of the research presented in this paper and those of the PrediHealth project (e.g., stratifying individuals into patients at risk or not). Specifically, we decided to keep all data related to a standard clinical and instrumental cardiological examination (without null values). Therefore, key echocardiographic measurements were retained, including left ventricular mass index (LVMI), ventricular dimensions, atrial size, posterior wall thickness, septum thickness, right ventricular diameter, and tricuspid annular systolic excursion (TAPSE). Electrical abnormalities on electrocardiographic evaluation, such as left and right bundle branch blocks (BBSx, BBDx), atrial fibrillation (FA), atrial flutter, and pacemaker presence (PM), were also included. Height and weight were removed since they are inherently represented by the BMI feature. In contrast, other variables such as sodium, potassium, uric acid, urea, free triiodothyronine (FT3), and free thyroxine (FT4) were excluded as they were considered secondary to the main goal of our research. Some other features were deliberately excluded due to the high percentage of missing values. Given the medical nature of the dataset, we opted not to impute missing values with synthetic or inferred data to preserve the integrity of the analysis and avoid introducing potential biases.
@@
Medication usage data, including beta-blockers, ACE inhibitors, and aldosterone antagonists, were incorporated to assess treatment effects. In addition, two derived characteristics were included: “lifespan,” representing the number of days between diagnosis and outcome, and “label,” a binary classification indicating whether a patient survived beyond the 440-day threshold. The features, grouped in “Clinical Features” and “Echocardiographic Features,” that serve as the foundation of our predictive model are detailed in Table I.
@@
NYHA New York Heart Association (NYHA) class (I-IV), measures heart failure severity Numeric
@@ III-B Predictive Model Workflow
Our predictive model is based on an ensemble learning approach, a modified stacking technique. Standard stacking combines multiple heterogeneous specialized models trained on the same feature set, our model employs domain-specific feature segregation to enhance predictive accuracy. Specifically, the input data are divided into two distinct feature groups: clinical features, including demographic data, patient history, laboratory results, and risk factors, and echocardiographic features. Each feature group is processed separately using two identical ML models, each trained exclusively on one of the feature sets. This separation ensures that each model specializes in learning patterns from its respective data type without interference from other features. In our implementation, both the specialized models and the meta-model are constructed using Random Forest algorithms. In detail, each specialized model is tuned using a five-fold cross-validation strategy, where the training data are partitioned into five subsets. In detail, for each specialized model, we employ a five-fold cross-validation strategy using GridSearchCV, which systematically tests various combinations of hyperparameters, such as the number of trees, maximum depth, and minimum samples required for a split, to determine the best configuration that minimizes prediction error. Once these models are optimized, a single train-test split (with 20% of the data reserved for testing) is utilized to evaluate the final performance. The outputs of the two optimized specialized models on the test set are then used as input features for the meta-model. Our predictive model workflow is inspired by the research presented by Pinem et al. [6], where the authors trained multiple models on the same set of features to predict heart failure risk. In that work, the authors used the output of the models as input for a Voting Classifier. In our case, the meta-model on top is also trained on the output of the specialized models, but the main difference is that each specialized model is trained using different features.
The outcome variable of the two models (i.e., the output of the two specialized models) is the estimated lifespan of a patient. The outputs of the two specialized models represent the input for a meta-model, which integrates their predictions to generate the final category for the patients. Specifically, the final output of the meta-model is expressed in terms of estimated lifespan. As mentioned before, the individuals for whom the model predicts a lifespan of less than 440 days are considered as patients “at risk,” while the other are considered as “not at risk.” In the context of the PrediHealt project, the identification of patients at HF risk allows them to participate in a telemonitoring program, with the objective of acting promptly to let them receive medical attention.
@@
Integrating two specialized models within a metamodel offers several potential advantages over a single ML model trained on all features. Specifically, these advantages are:
@@
    Improved Interpretability and Modularity. By separating clinical and echocardiographic features, the contribution of each data type can be analyzed independently. This modular approach allows for better insight into how different variables influence predictions.
@@
    Reduced Risk of Overfitting. A single model trained on all features may overfit due to redundant or spurious correlations. By using specialized models for each group of features (clinical and echocardiographic), we reduce the risk of one type of feature dominating the learning process, improving generalization.
@@
    Increased Robustness. This structure makes the system less sensitive to missing data. If a patient lacks echocardiographic measurements, the clinical model can still contribute to predictions, preventing the complete failure of the model. The metamodel also balances predictions, reducing the impact of errors from individual models.
@@
  5. 5.
    Greater Adaptability and Scalability. If new predictive techniques emerge for clinical or echocardiographic data, only the respective model needs to be updated, rather than retraining an entire model on all the features. This characteristic allows for continuous improvement of the predictions while maintaining the same structure of the predictive framework.

### v3

# Machine Learning Solutions Integrated in an IoT Healthcare Platform for Heart Failure Risk Stratification

@@ Abstract
This paper presents a predictive model founded on Machine Learning (ML) techniques to identify patients at Heart Failure (HF) risk. This model is an ensemble learning approach, a modified stacking technique, that uses three ML models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these three models. We assessed the model on a real dataset, and our results suggest that it performs well in the stratification of patients at HF risk. Specifically, we obtained high sensitivity (91%) and good accuracy (78%). To better study the value and potentiality of our model, we also contrasted its results with those obtained by using several baseline ML models. The preliminary results indicate that our predictive model outperforms these baselines that flatly consider features, i.e., not grouping features in clinical and echocardiographic features. The proposed model is a part of the PrediHealth research project. This project integrates telemedicine, mobile health solutions, and predictive analytics into an IoT healthcare platform to manage patients at HF risk.
@@ I Introduction
The human heart works like a pump, sending blood and oxygen around the body. A major cause of morbidity and mortality worldwide is chronic Heart Failure (HF), which is defined by the heart’s inability to pump blood efficiently. Ageing of population and lifestyle risk factors such as smoking, poor diet, and high blood pressure are closely associated with the increase in the incidence of heart disease. Current techniques often fail to provide prompt and precise identification of heart disease, even with advances in medical diagnostics [2, 3]. It is also worth mentioning that cardiovascular diseases are associated with the highest number of deaths among Non-Communicable Diseases, amounting to 17.9 million deaths annually, followed by cancers (9.3 million), chronic respiratory diseases (4.1 million), and diabetes (2 million, including kidney disease linked to diabetes) [4].
@@
The PrediHealth project aims to address the challenges above by integrating telemedicine, mobile health (m-health) solutions, and predictive analytics into an IoT healthcare platform. Specifically, it leverages an interoperable IoT web-based application, a telemonitoring kit equipped with medical devices and environmental sensors, and ML/AI-driven predictive models to support clinical decisions [5]. As it is easy to follow, at the core of the PrediHealth project lies the predictive model–designed for the stratification of patients at HF risk–that is presented in this paper for the first time. This model is based on an ensemble learning approach–a modified stacking technique–that uses three ML models leveraging clinical and echocardiographic features and then a meta-model to combine the predictions of these three bottom models. In this paper, we detail our model and its initial empirical assessment, as well as the possible implications for our results inside and outside the goal of the PrediHealth project. We assessed the validity of the model by conducting an initial empirical study. The obtained results suggest that the model performs pretty well. We obtained high sensitivity (91%), ensuring that such a model nearly identifies all HF patients. As for accuracy, we obtained 78%. It is acceptable in the PrediHealth project, given the priority of identifying patients at risk of HF to be then included in a telemonitoring program, to act promptly to let them receive medical attention by using the IoT healthcare platform developed in the project. We also compared our model against three baseline models (e.g., Decision Tree and Support Vector Regression). Our initial outcomes suggest that the use of our model outperforms the baseline models (e.g., in terms of the accuracy and the sensitivity).
@@
Paper Structure. In Section II, we introduce the PrediHealth project and present related research and background information, while the design of the predictive model is shown in Section III. The design of the study to empirically assess our predictive model and the results of this assessment are reported in Section IV and Section V, respectively. In this latter section, we also discuss the results and delineate possible implications for our research. We conclude the paper in Section VI with final remarks and future work.
@@ II Background and Related Work
In this section, we first provide information on the PrediHealth project. It is the project for which the predictive model presented in this paper has been developed. Then we continue motivating our work and discussing related research. We conclude with a brief introduction to the ML models used in our work.
@@ II-A The PrediHealth Project
We restricted ourselves to a few aspects of the project for space reasons and scant relevance to the overarching goals of the research presented in the paper. Readers interested in further details can refer to [5].
@@ II-B Predictive Models in Digital Health
Other researchers have explored ML-driven models for HF prognosis, utilizing algorithms like logistic regression, decision trees, random forests, support vector machines (SVMs), and ensemble classifiers. For instance, Pinem et al. [7] demonstrated that a Voting Classifier combining Logistic Regression and SVM achieved the highest accuracy (88.04%) and ROC_AUC score (88%), with blood pressure and cholesterol emerging as key predictive factors. Similarly, Guo et al. [8] employed ML techniques to predict HF outcomes using EHR data, incorporating medical notes, laboratory results, and imaging data to achieve expert-level diagnostic accuracy. These studies underscore the role of ML in early HF detection, risk assessment, and personalized intervention strategies. Recently, Visco et al. [9] developed a predictive model for cardiovascular diagnostics. Their study identified correlations between worsening HF and three key parameters: Creatinine, systolic pulmonary artery pressure (sPAP), and coronary artery disease(CAD). To enhance interpretability, they introduced a GP-based classifier and extracted from it a formula that achieved almost 97% accuracy, outperforming traditional ML models. However, the authors themselves identified how the use of only four features, despite the high accuracy, could leave out other parameters that could be important in the final prediction from the doctor’s final evaluation.
@@
The research presented in this paper aims to bridge these gaps by developing an ML-based predictive model that integrates diverse patient datasets, including HFpEF cases, to enhance generalizability. The proposed model not only refines risk stratification but also enhances clinical applicability by ensuring interpretability and reliability across heterogeneous patient populations. In addition, our predictive model is integrated within advanced solutions in the context of CPS/IoT fields [^2] While CPS encompasses a broader range of networked systems, IoT specifically focuses on internet-enabled connectivity. by integrating state-of-the-art technologies such as telemonitoring and m-health to facilitate the integrated hospital-territory-home management of HF patients. Finally, integrating three models within a metamodel offers several advantages over a single ML model trained on all features. We detail these advantages in Section III-B.
II-C Used ML-Models
We used two machine learning models: the Logistic Regression Classifier and the Random Forest Classifier. The former is a linear model widely applied in binary classification problems [13]. It applies a weighted sum of the input features and uses the logistic function to estimate the probability of a given class. Due to its simplicity, it is computationally efficient and generally less prone to overfitting, especially in low-dimensional spaces. These characteristics, along with the inherent interpretability of the model’s results (i.e., how easily a person can trace and explain the output based on the input features), make the Logistic Regression Classifier particularly valuable in clinical settings [14], where understanding model decisions is of central importance. In contrast, the Random Forest Classifier is a non-linear, ensemble-based method that constructs a collection of decision trees during the training phase and aggregates their outputs to form a final prediction [15]. This approach enables the model to capture intricate patterns and interactions across features, particularly in heterogeneous or high-dimensional datasets. While Random Forests are generally robust and versatile, they tend to lack the transparency of linear models, which may be a drawback in domains where interpretability is a key requirement.
@@ III Design of the Predictive Model
In this section, we first introduce the dataset employed for the training and the testing of our model, and then we describe it in detail.
@@ III-A Dataset
The used dataset contains real medical data. That is, it comprises clinical records of patients diagnosed with heart failure, collected as part of a structured medical registry. It encompasses a diverse set of 1040 patient records, each characterized by unique ID and 65 features that provide a comprehensive view of patient demographics, clinical history, treatment details, and outcomes. The dataset includes key variables such as age, height, weight, and body mass index (BMI), along with clinical indicators like ejection fraction (EF), New York Heart Association (NYHA) classification, and specific diagnoses, including ischemic, hypertensive, and dilated cardiomyopathy. Additionally, dataset documents prescribed medications, including ACE inhibitors, Beta-blockers, and aldosterone antagonists, offering insights into pharmacological management. Longitudinal follow-up data track patient progress, recording mortality outcomes such as death, heart failure exacerbation, and interventions like heart transplants or ventricular assist device (VAD) implantation. The dataset also contains, e.g., other bio-chemical pieces of information like Na+ (Sodium) or Uric Acid.
III-A1 Features
@@
We extracted 33 out of 65 available features in the dataset, including clinical and echocardiographic parameters relevant to cardiac evaluation based on the experience of some of the authors in the HF field and our knowledge of the state of the art (e.g., [9]). The chosen characteristics encompass demographic attributes such as age, BMI, and sex, along with essential clinical indicators such as ejection fraction (EF), NYHA classification, NT-proBNP, creatinine, glucose, and hemoglobin (Hb). Additionally, the dataset contains information on comorbidities such as hypertension, diabetes, dyslipidemia, and respiratory diseases. The selection of features was based on the experience some of the authors gained in the field of HF, taking into account the purposes of the research presented in this paper and those of the PrediHealth project (e.g., stratifying individuals into patients at risk or not). Specifically, we decided to keep all data related to a standard clinical and instrumental cardiological examination (without null values). Therefore, key echocardiographic measurements were retained, including left ventricular mass index (LVMI), ventricular dimensions, atrial size, posterior wall thickness, septum thickness, right ventricular diameter, and tricuspid annular systolic excursion (TAPSE). Electrical abnormalities on electrocardiographic evaluation, such as left and right bundle branch blocks (BBSx, BBDx), atrial fibrillation (FA), atrial flutter, and pacemaker presence (PM), were also included. Height and weight were removed since they are inherently represented by the BMI feature. In contrast, other variables such as sodium, potassium, uric acid, urea, free triiodothyronine (FT3), and free thyroxine (FT4) were excluded as they were considered secondary to the main goal of our research. It is important to note that the Diagnosis feature was divided into two categorical features to enhance the granularity of the information and allow for a clearer differentiation between the identified diagnoses. Some other features were deliberately excluded due to the high percentage of missing values. Given the medical nature of the dataset, we opted not to impute missing values with synthetic or inferred data to preserve the integrity of the analysis and avoid introducing potential biases.
@@
Medication usage data, including beta-blockers, ACE inhibitors, and aldosterone antagonists, were incorporated to assess treatment effects. In addition, two derived characteristics were included: “lifespan,” representing the number of days between diagnosis and outcome, and “label.” The features, grouped in “Clinical Features” and “Echocardiographic Features,” that serve as the foundation of our predictive model are detailed in Table I.
@@
NYHA New York Heart Association (NYHA) class (I-IV), measures heart failure severity Categorical
@@
III-A2 Pre-processing
Before using the dataset, we preprocessed it to ensure data consistency and reliability. Therefore, the goal of pre-processing was to include the highest number of data and have, at the same time, a balance between patients at HF risk and not. The following pre-processing steps were performed:
  • Date values were converted into a standardized date-time format, and a new feature, “lifespan,” was introduced to represent the number of days between the initial characterization date and the recorded outcome date (“death”).
  • We hold patients for whom we had information on their death or the follow-up period was larger or equal to three years. At the end of this step, 464 patients remained in the dataset.
  • Patients with at least one missing value in the features used by our model were removed from the dataset. At the end of this step, we removed 99 patients from the dataset, and then 365 patients remained.
  • Patients with inconsistent values or placeholder values in terms of classification (intermediate values) have been removed. At the end of this step, we removed 8 patients from the dataset, and we reached the final dimension of the dataset of 357 patients
  • As for the outcome variable (or also dependent variable), we used a binary classification for the values in the dataset. We employed a threshold-based labeling approach, using a three-year threshold to classify patients into two groups: “HF at risk” and “not at HF risk.” The choice of this threshold is justified by technical reasons and clinical bases. Indeed, from a technical point of view, this threshold divides the dataset almost perfectly; in fact, we have 45% of samples with the label “at risk” and 55% of samples with the label “not at risk.” This also allowed the predictive model to exhibit good prediction performance of the outcome values. From a clinical perspective, the chosen threshold aligns with a time frame during which close monitoring and proactive management are essential to mitigate adverse outcomes in patients with HF.
@@ III-B Predictive Model Workflow
Our predictive model (also meta-model or simply model from here onwards) is based on an ensemble learning approach, a modified stacking technique, to classify patients at HF risk or not. Standard stacking combines multiple heterogeneous models trained on the same set of features, our approach leverages domain-specific feature segregation to enhance predictive accuracy. The input data are divided into two distinct feature groups: clinical features, including demographic data, patient history, laboratory results, and risk factors, and echocardiographic features. Each feature group is processed separately using two identical ML models (i.e., also clinical and echocardiographic specialized models from here onwards), each trained exclusively on one of the feature sets. This separation ensures that each model specializes in learning patterns from its respective data type without interference from other features. Both the specialized models are implemented using Logistic Regression Classifiers. In detail, we employ a five-fold cross-validation strategy using GridSearchCV, which systematically tests various combinations of hyperparameters for a split to determine the best configuration that maximizes the accuracy. In addition to these two models, we employ a third model trained on the complete set of features (i.e., the merge between clinical and echocardiographic features). Indeed, we used a Random Forest Classifier to capture complex and non-linear relationships across the feature space that could be difficult to detect for Logistic regression classifiers. Using all the features together also allows for capturing interactions and patterns that span across different data types, potentially uncovering valuable insights that might be lost when the two sets of features are treated separately. A single train-test split (with 20% of the data reserved for testing) is exploited to evaluate the performance of these three models. The selection of the three models used was informed by preliminary empirical evidence that indicated their effectiveness alone and in the combined use in the meta-model, given the specific domain of our research and the peculiarities of our dataset, .
The outputs of the three models described above, along with their confidence in the prediction, are used as inputs for our meta-model (i.e., a classifier), which is implemented as a Logistic Regression Classifier responsible for aggregating the predictions of these three models into a single prediction. The idea underlying our predictive model is inspired by the research presented by Pinem et al. [7], where the authors trained multiple models on the same set of features to predict heart failure risk. In that work, the authors used the output of the models as input for a Voting Classifier. In contrast, our method introduces two key differences: (i) the use of specialized models trained on disjoint subsets of features (supported by a model trained on the complete feature set) and (ii) the use of a learnable meta-model (rather than a static voting mechanism) for the prediction.
@@
Integrating a set of models within a meta-model offers several potential advantages over a single ML model:
@@
    Interpretability and Modularity. By separating clinical and echocardiographic features, the contribution of each data type can be analyzed independently. Additionally, a third model trained on all features captures joint relationships across the data. The study of these models separately also allows for a clearer interpretation of which type of information most influences the final predictions. We studied these aspects in our empirical assessment of the proposed solutions.
@@
    Risk of Overfitting. A single model trained on all features may overfit due to redundant or spurious correlations. By using specialized models for each group of features (clinical and echocardiographic), we reduce the risk of one type of feature dominating the learning process, improving generalization.
@@
    Robustness. This structure makes the system less sensitive to missing data. If a patient lacks echocardiographic measurements, the clinical model can still contribute to predictions, preventing the complete failure of the model. The metamodel also balances predictions, reducing the impact of errors from individual models.